OpenStudy (rania):

If 3 J of work are needed to stretch a spring from 11 cm to 13 cm and 5 J are needed to stretch it from 13 cm to 15 cm, what is the natural length of the spring?

OpenStudy (rcp031):

what 3 J & 5 J mean?

OpenStudy (mathmate):

J=joule for more information, read https://en.wikipedia.org/wiki/Joule Here the two unknowns are 1. spring constant, k (stiffness of spring) 2. natural length, L The standard relation between stretching and work done (in joules) is W(x1,x2)=(k/2)(L+x2)^2-k(L+x1)^2 where the string is stretched from L+x1 to L+x2. L=natural of spring (m) W=work done (J) k=spring constant (J/m)  Two equations can be formed, one for each case/observation: 3=(k/2)((L+0.13)^2-(L+0.11)^2) 5=(k/2)((L+0.15)^2-(L+0.13)^2) Solve the system of equations and find k and L.